# Composing Codensity Bisimulations

### v1

@@ 6.1 Transferring Liftable Modalities
We next show that the transfer of liftable modalities from one ${\mathbf{CLat}_{\sqcap}}$-fibration to another. We first see the interaction between $L/R$ and the $\mathrm{Sp}$.
@@
When $F=Id$, these properties are working behind Komorida et al.’s argument of transferring codensity bisimilarities [23]. Let $(Id,G):p\to q$ be a 1-cell in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ such that $G$ is full, fibered and preserving fiberwise meets; such a 1-cell is called a transfer situation in [23]. Then the 1-cell commutes with codensity liftings, that is, $G\circ[F,\tau]^{\mathbf{\Omega}}=[F,\tau]^{G\circ\mathbf{\Omega}}\circ G$. From this commutativity, together with the adjoint lifting theorem of Hermida and Jacobs [17, Theorem 2.14] (see also [29, 23, 32] for relevant results), the preservation of the codensity bisimilarity is obtained:
$G(\nu(c^{*}\circ[F,\tau]^{\mathbf{\Omega}}))=\nu(c^{*}\circ[F,\tau]^{G\circ\mathbf{\Omega}})\quad(c:X\to F(X)).$
@@
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $(T,\lambda)$ be an $N$-ary composition operation for $F$-coalgebras. Let $(Id,G):p\to(q:\mathbb{F}\to\mathbb{B})$ be a 1-cell in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ such that $G$ is full, fibered and preserving fibered meets. If a modality $\sigma$ lifts $\lambda$ along $p$, then in the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $((\Omega,G\circ\mathbf{\Omega}):Id_{A}\to(q:\mathbb{F}\to\mathbb{B}))$, $\sigma$ lifts $\lambda$ along $q$
@@
Consider the ${\mathbf{CLat}_{\sqcap}}$-fibration $U_{\mathbf{ERel}}\colon\mathbf{ERel}\to\mathbf{Set}$ and the inclusion functor $i\colon\mathbf{EqRel}\hookrightarrow\mathbf{ERel}$. Then $(U_{\mathbf{EqRel}},U_{\mathbf{ERel}},i)$ is a transfer situation. In Example 6.5, we showed that $\sigma_{\land}$ lifts $\lambda^{\mathcal{P}}$ along $U_{\mathbf{EqRel}}$. Proposition 6.12 induces that $\sigma_{\land}$ lifts $\lambda^{\mathcal{P}}$ also along $U_{\mathbf{ERel}}$.
@@ 7.1 Bisimilarity Pseudometric for Deterministic Automata
We next visit bisimilarity pseudometric for deterministic automata [2, 5], which is a quantitative extension of language equivalence. We fix a weight $w\in\mathbb{I}$. We regard the interval $\mathbb{I}$ as a quantitative extension of binary truth values via the cast function $i:2\to\mathbb{I}$ defined by $i(\mathrm{true})=0$ and $i(\mathrm{false})=1$.
@@
We use the same coalgebraic formulation of deterministic automata as ${F_{\mathrm{da}}}$-coalgebras as Example 6.7. The difference from that section is that we employ the category of pseudometric spaces for modeling bisimilarity pseudometric. We adopt the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega}):Id_{\Sigma\uplus\{\epsilon\}}\to U_{\mathbf{PMet}}$ where $\mathbf{\Omega}$ constantly returns the Euclidean space $(\mathbb{I},d_{\mathbb{I}})$ over the $[0,1]$-interval. We pair it with the behavior functor ${F_{\mathrm{da}}}$ for deterministic automata and the following modality $\tau_{a}\colon{F_{\mathrm{da}}}(\mathbb{I})\to\mathbb{I}$ to obtain codensity bisimulation data:
@@
The $[{F_{\mathrm{da}}},\tau]^{\mathbf{\Omega}}$-bisimilarity on a ${F_{\mathrm{da}}}$-coalgebra $c:X\to{F_{\mathrm{da}}}(X)$ is a quantitative extension of the language equivalence on $X$. It maps $(x,x^{\prime})\in X^{2}$ to 00 if the languages of $x,x^{\prime}$ are the same, and $w^{n}$ otherwise where $n$ is the minimum length of a word that is accepted from one and not from the other. This notion corresponds to language equivalence if $w=1$.
@@ 7.2 Similarity Pseudometric for Deterministic Automata
Let us consider similarity for the previous subsection. We get a similarity framework just by relaxing the symmetry condition of $\mathbf{PMet}$ and taking an asymmetric truth-value domain instead of the Euclidean distance $d_{\mathbb{I}}$. For omitted proofs, see Appendix A.3.2.
@@
Let $\mathbf{LMet}$ be the category of Lawvere metric spaces (meaning asymmetric pseudometric spaces) and non-expansive maps. The forgetful functor $U_{\mathbf{LMet}}\colon\mathbf{LMet}\to\mathbf{Set}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration like $U_{\mathbf{PMet}}$. We adopt the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega}):Id_{A}\to U_{\mathbf{LMet}}$ such that $\mathbf{\Omega}$ constantly returns the space over the interval with the Lawvere metric $d_{\mathbb{I}}^{\mathrm{as}}(x,y)\coloneqq\max(0,y-x)$. We pair it with the behavior functor ${F_{\mathrm{da}}}$ for deterministic automata and the same modality $\tau$ in §7.1 to form codensity bisimulation data.
@@
recall that $i:2\to\mathbb{I}$ is the cast function (§7.1). Note that $d_{\mathbb{I}}^{\mathrm{as}}$ maps $(\mathrm{true},\mathrm{false})$ to $1$ and the others to 00. In §7.1 we used the euclidean distance $d_{\mathbb{I}}$ instead of $d_{\mathbb{I}}^{\mathrm{as}}$, which also maps $(\mathrm{false},\mathrm{true})$ to $1$. It makes a difference in the bisimilarity as below.
The $[{F_{\mathrm{da}}},\tau]^{\mathbf{\Omega}}$-bisimilarity on a ${F_{\mathrm{da}}}$-coalgebra $c:X\to{F_{\mathrm{da}}}(X)$ is a similarity pseudometric on $X$ for the deterministic automaton $c$. It maps $(x,x^{\prime})\in X^{2}$ to 00 if the language of $x$ is included in that of $x^{\prime}$, and $w^{n}$ otherwise where $n$ is the minimum length of a word that is accepted from $x$ and not from $x^{\prime}$. A bisimialrity pseudometric in §7.1 detects words that are accepted from $x^{\prime}$ and not from $x$ while a similarity pseudometric doesn’t.
@@ 7.3 Bisimulation Metric for Markov Decision Processes
We first set-up coalgebraic formulation of Markov decision processes (MDPs) and bisimulation metric. Let us write ${\mathcal{D}}\colon\mathbf{Set}\to\mathbf{Set}$ for the probability distribution functor which maps $X$ to the set of distributions on $X$. Then MDPs are naturally modeled by ${\mathcal{P}}\circ{\mathcal{D}}$-coalgebras $c:X\to{\mathcal{P}}{\mathcal{D}}(X)$. Hereafter we omit the composition between ${\mathcal{P}}$ and ${\mathcal{D}}$.
@@
The first codensity bisimulation data is $((\mathbb{I},d_{\mathbb{I}}):Id_{1}\to U_{\mathbf{PMet}},{\mathcal{D}},e)$ where $e$ is the expectation function $e\colon{\mathcal{D}}\mathbb{I}\to\mathbb{I}$. The codensity lifting $[{\mathcal{D}},e]^{d_{\mathbb{I}}}\colon\mathbf{PMet}\to\mathbf{PMet}$ maps $d\in\mathbf{PMet}_{X}$ to the Kantorovich pseudometric $\mathcal{K}(d)$ on ${\mathcal{D}}X$:
@@
The second codensity bisimulation data is $((\mathbb{I},d_{\mathbb{I}}):Id_{1}\to U_{\mathbf{PMet}},{\mathcal{P}},\mathrm{inf})$. The codensity lifting $[{\mathcal{P}},\mathrm{inf}]^{d_{\mathbb{I}}}\colon\mathbf{PMet}\to\mathbf{PMet}$ mapping $d\in\mathbf{PMet}_{X}$ to the Hausdorff $\mathcal{H}(d)$ pseudometric on ${\mathcal{P}}X$:
@@
Then a $[{\mathcal{P}},\mathrm{inf}]^{d_{\mathbb{I}}}\circ[{\mathcal{D}},e]^{d_{\mathbb{I}}}$-bisimulation on $c$ is called a bisimulation metric, and the bisimilarity is called a bisimilarity metric [7]. It maps a pair of states $(x,y)$ to a distance that measures quantitative analogue to behavioral equivalences.
@@
We have seen the binary one-step composition (×$(\times,\lambda^{{\mathcal{P}}})$ for ${\mathcal{P}}$ in Example 6.5. We next introduce the one for ${\mathcal{D}}$ by
@@
$\lambda^{{{\mathcal{P}}{\mathcal{D}}}}\triangleq({\mathcal{P}}\circ\lambda^{\mathcal{D}})\bullet(\lambda^{\mathcal{P}}\circ{\mathcal{D}}^{2}).$
@@
As a consequence of the proposition above, given two MDPs $c_{i}:X_{i}\to{{\mathcal{P}}{\mathcal{D}}}(X_{i})$ over at most countable states $X_{i}$ ($i=1,2$), if $d_{i}$ is a bisimulation metric for $c_{i}$, then the mapping
$d((x,y),(x^{\prime},y^{\prime}))\triangleq{\sigma_{\oplus}}(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))$
@@
The second statement comes from the following lemma, which is a direct adaptation of [13, Theorem 2.20] in the context of pseudometric spaces. It can be proved in almost the same way as the original theorem; here we rely on the Kantorovich-Rubinstein duality theorem on at most countable sets [35].
@@
Let $X,Y$ be at most countable sets, $(X,d_{1}),(Y,d_{2})\in\mathbf{PMet}$, and $\sigma:\mathbb{I}^{2}\to\mathbb{I}$ be a concave function satisfying (6). Then for each $\mu_{1},\mu_{1}^{\prime}\in{\mathcal{D}}X$ and $\mu_{2},\mu_{2}^{\prime}\in{\mathcal{D}}Y$,
@@
We didn’t apply Theorem 6.2 to this case because as of now, it remains unsolved that the modality ${\sigma_{\oplus}}$ satisfies the second condition in Theorem 6.2 while it satisfies the first one.
@@ 8 Composing Codensity Games
We introduce a composition of codensity games, and show that it preserves game invariants. As a consequence, we provide an alternative proof of the preservation of bisimilarities (the inequality shown in Corollary 3.10) under the sufficient condition in Theorem 6.2. Equivalently, the preservation of bisimilarities is restated as follows:
$\forall P_{1},\cdots,P_{N}~{}.~{}$ $\textstyle\bigwedge_{i\in N}\left(P_{i}\sqsubseteq\nu(c_{i}^{*}\circ[F,\tau]^{\mathbf{\Omega}})\right)$
$\implies[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})\sqsubseteq\nu(T_{\lambda}(c_{1},\cdots,c_{N})^{*}\circ[F,\tau]^{\mathbf{\Omega}}).$
We call $P_{i}$ a witness of bisimilarity, and call the above formula the compositionality of witnesses.
Definition 8.1 (witness).
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $c\colon X\to F(X)$ be an $F$-coalgebra. An object $P\in\mathbb{E}_{X}$ is a witness of the codensity bisimilarity on $c$ if $P\sqsubseteq\nu\bigl{(}c^{*}\circ[F,\tau]^{\mathbf{\Omega}}\bigr{)}$.

### v2

@@ 6.2 Transferring Liftable Modalities
We next study the transfer of liftable modalities from one ${\mathbf{CLat}_{\sqcap}}$-fibration to another. We first see the interaction between $L/R$ and the $\mathrm{Sp}$.
@@
When $F=Id$, these properties are working behind Komorida et al.’s argument of transferring codensity bisimilarities [23]. Let $(Id,G)\colon p\to q$ be a 1-cell in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ such that $G$ is full, fibered and preserving fiberwise meets; such a 1-cell is called a transfer situation in [23]. Then the 1-cell commutes with codensity liftings, that is, $G\circ[F,\tau]^{\mathbf{\Omega}}=[F,\tau]^{G\circ\mathbf{\Omega}}\circ G$. From this commutativity, together with the adjoint lifting theorem of Hermida and Jacobs [17, Theorem 2.14] (see also [29, 23, 32] for relevant results), the preservation of codensity bisimilarity is obtained:
$G(\nu(c^{*}\circ[F,\tau]^{\mathbf{\Omega}}))=\nu(c^{*}\circ[F,\tau]^{G\circ\mathbf{\Omega}})\quad(c\colon X\to F(X)).$
@@
Let $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $(T,\lambda)$ be an $N$-ary composition operation for $F$-coalgebras. Let $(Id,G)\colon p\to(q\colon\mathbb{F}\to\mathbb{B})$ be a 1-cell in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ such that $G$ is full, fibered and preserving fibered meets. If a modality $\sigma$ lifts $\lambda$ along $p$, then $\sigma$ lifts $\lambda$ also along $q$ in the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $((\Omega,G\circ\mathbf{\Omega})\colon Id_{A}\to(q\colon\mathbb{F}\to\mathbb{B}))$.
@@
Consider the ${\mathbf{CLat}_{\sqcap}}$-fibration $U_{\mathbf{ERel}}\colon\mathbf{ERel}\to\mathbf{Set}$ and the inclusion functor $i\colon\mathbf{EqRel}\hookrightarrow\mathbf{ERel}$. Then $(U_{\mathbf{EqRel}},U_{\mathbf{ERel}},i)$ is a transfer situation. In §6.1.1, we showed that $\sigma_{\land}$ lifts $\lambda^{\mathcal{P}}$ along $U_{\mathbf{EqRel}}$. Proposition 6.10 induces that $\sigma_{\land}$ lifts $\lambda^{\mathcal{P}}$ also along $U_{\mathbf{ERel}}$.
@@ 7.1 Bisimilarity Pseudometric for Deterministic Automata
We next study a bisimilarity pseudometric for deterministic automata [2, 5], which is a quantitative extension of language equivalence. We fix a weight $w\in\mathbb{I}$ where $\mathbb{I}$ is the interval $[0,1]$. We regard the interval $\mathbb{I}$ as a quantitative extension of binary truth values via the cast function $i\colon 2\to\mathbb{I}$ defined by $i(\mathrm{true})=0$ and $i(\mathrm{false})=1$.
@@
We use the same coalgebraic formulation of deterministic automata as ${F_{\mathrm{da}}}$-coalgebras as §6.1.2 where the functor ${F_{\mathrm{da}}}$ is introduced in Example 4.3. The difference from that example is that we employ the category of pseudometric spaces for modeling bisimilarity pseudometrics. We adopt the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega})\colon Id_{\Sigma\uplus\{\epsilon\}}\to U_{\mathbf{PMet}}$ where $\mathbf{\Omega}$ constantly returns the Euclidean space $(\mathbb{I},d_{\mathbb{I}})$ over the $[0,1]$-interval. We pair it with the behavior functor ${F_{\mathrm{da}}}$ for deterministic automata and the following modality $\tau_{a}\colon{F_{\mathrm{da}}}(\mathbb{I})\to\mathbb{I}$ to obtain codensity bisimulation data:
@@
Now, $[{F_{\mathrm{da}}},\tau]^{\mathbf{\Omega}}$-bisimilarity on an ${F_{\mathrm{da}}}$-coalgebra $c\colon X\to{F_{\mathrm{da}}}(X)$ is a quantitative extension of language equivalence on $X$. It maps $(x,x^{\prime})\in X^{2}$ to 00 if the languages of $x,x^{\prime}$ are the same, and $w^{n}$ otherwise where $n$ is the minimum length of a word that is accepted from one and not from the other. This notion corresponds to language equivalence if $w=1$.
@@ 7.2 Similarity Pseudometric for Deterministic Automata
Let us consider similarity on deterministic automata, as a variation of the example in the previous subsection.
We get a similarity framework just by relaxing the symmetry condition of $\mathbf{PMet}$ and taking an asymmetric truth-value domain instead of the Euclidean distance $d_{\mathbb{I}}$. For omitted proofs, see Appendix A.3.2.
@@
Let $\mathbf{LMet}$ be the category of Lawvere metric spaces (meaning asymmetric pseudometric spaces) and non-expansive maps. The forgetful functor $U_{\mathbf{LMet}}\colon\mathbf{LMet}\to\mathbf{Set}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration like $U_{\mathbf{PMet}}$. We adopt the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega})\colon Id_{A}\to U_{\mathbf{LMet}}$ such that $\mathbf{\Omega}$ constantly returns the space over the interval with the Lawvere metric $d_{\mathbb{I}}^{\mathrm{as}}(x,y)\coloneqq\max(0,y-x)$. We pair it with the behavior functor ${F_{\mathrm{da}}}$ for deterministic automata and the same modality $\tau$ in §7.1 to form codensity bisimulation data.
@@
recall that $i\colon 2\to\mathbb{I}$ is the cast function (§7.1). Note that $d_{\mathbb{I}}^{\mathrm{as}}$ maps $(\mathrm{true},\mathrm{false})$ to $1$ and the others to 00. In §7.1 we used the Euclidean distance $d_{\mathbb{I}}$ instead of $d_{\mathbb{I}}^{\mathrm{as}}$, which also maps $(\mathrm{false},\mathrm{true})$ to $1$.
The choice of the Euclidean distance makes a difference in the notion of bisimilarity. Now $[{F_{\mathrm{da}}},\tau]^{\mathbf{\Omega}}$-bisimilarity on an ${F_{\mathrm{da}}}$-coalgebra $c\colon X\to{F_{\mathrm{da}}}(X)$ is a similarity pseudometric on $X$ for the deterministic automaton $c$. It maps $(x,x^{\prime})\in X^{2}$ to 00 if the language of $x$ is included in that of $x^{\prime}$, and $w^{n}$ otherwise where $n$ is the minimum length of a word that is accepted from $x$ and not from $x^{\prime}$. A bisimialrity pseudometric in §7.1 detects words that are accepted from $x^{\prime}$ and not from $x$ while a similarity pseudometric doesn’t.
@@ 7.3 Bisimulation Metric for Markov Decision Processes
We first set-up coalgebraic formulation of Markov decision processes (MDPs) and bisimulation metric. Let us write ${\mathcal{D}}\colon\mathbf{Set}\to\mathbf{Set}$ for the probability distribution functor which maps $X$ to the set of distributions on $X$. Then MDPs are naturally modeled by ${\mathcal{P}}\circ{\mathcal{D}}$-coalgebras $c\colon X\to{\mathcal{P}}{\mathcal{D}}(X)$. Hereafter we omit the composition between ${\mathcal{P}}$ and ${\mathcal{D}}$.
@@
The first codensity bisimulation data is $((\mathbb{I},d_{\mathbb{I}})\colon Id_{1}\to U_{\mathbf{PMet}},{\mathcal{D}},e)$ where $e$ is the expectation function $e\colon{\mathcal{D}}\mathbb{I}\to\mathbb{I}$. The codensity lifting $[{\mathcal{D}},e]^{d_{\mathbb{I}}}\colon\mathbf{PMet}\to\mathbf{PMet}$ maps $d\in\mathbf{PMet}_{X}$ to the Kantorovich pseudometric $\mathcal{K}(d)$ on ${\mathcal{D}}X$:
@@
The second codensity bisimulation data is $((\mathbb{I},d_{\mathbb{I}})\colon Id_{1}\to U_{\mathbf{PMet}},{\mathcal{P}},\mathrm{inf})$. The codensity lifting $[{\mathcal{P}},\mathrm{inf}]^{d_{\mathbb{I}}}\colon\mathbf{PMet}\to\mathbf{PMet}$ mapping $d\in\mathbf{PMet}_{X}$ to the Hausdorff $\mathcal{H}(d)$ pseudometric on ${\mathcal{P}}X$:
@@
Then a $[{\mathcal{P}},\mathrm{inf}]^{d_{\mathbb{I}}}\circ[{\mathcal{D}},e]^{d_{\mathbb{I}}}$-bisimulation on $c$ is called a bisimulation metric, and the bisimilarity is called the bisimilarity metric [7]. It maps a pair of states $(x,y)$ to a distance that measures quantitative analogue to behavioral equivalences.
@@
We have seen the binary one-step composition (×$(\times,\lambda^{{\mathcal{P}}})$ for ${\mathcal{P}}$ in §6.1.1. We next introduce the one for ${\mathcal{D}}$ by
@@
$\lambda^{{{\mathcal{P}}{\mathcal{D}}}}\coloneq({\mathcal{P}}\circ\lambda^{\mathcal{D}})\bullet(\lambda^{\mathcal{P}}\circ{\mathcal{D}}^{2}).$
@@
As a consequence of the proposition above, given two MDPs $c_{i}\colon X_{i}\to{{\mathcal{P}}{\mathcal{D}}}(X_{i})$ over at most countable states $X_{i}$ ($i=1,2$), if $d_{i}$ is a bisimulation metric for $c_{i}$, then the mapping
$d((x,y),(x^{\prime},y^{\prime}))\coloneq{\sigma_{\oplus}}(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))$
@@
The second statement comes from the following lemma, which is a direct adaptation of [12, Theorem 2.20] in the context of pseudometric spaces. It can be proved in almost the same way as the original theorem; here we rely on the Kantorovich-Rubinstein duality theorem on at most countable sets [36].
@@
Let $X,Y$ be at most countable sets, $(X,d_{1}),(Y,d_{2})\in\mathbf{PMet}$, and $\sigma\colon\mathbb{I}^{2}\to\mathbb{I}$ be a concave function satisfying (6). Then for each $\mu_{1},\mu_{1}^{\prime}\in{\mathcal{D}}X$ and $\mu_{2},\mu_{2}^{\prime}\in{\mathcal{D}}Y$,
@@
We did not apply Theorem 6.2 to this case because as of now, it remains unsolved whether the modality ${\sigma_{\oplus}}$ satisfies the second condition in Theorem 6.2. We note that the modality satisfies the first condition.
@@ 8 Composing Codensity Games
We introduce a composition of codensity games, and show that it preserves game invariants. As a consequence, we prove the preservation of bisimilarities (the inequality shown in Corollary 3.10) under the sufficient condition in Theorem 6.2 by composition of games.